OpenStudy (anonymous):

Use the x-intercepts to find the intervals on which the graph of f is above and below the x-axis. f(x) = (x + 5)^2

OpenStudy (anonymous):

I think you factor then graph y=X^2+10x+25 then graph

OpenStudy (anonymous):

\[(- \infty,5) and (5,\infty)\]

OpenStudy (anonymous):

nope oops\[(-\infty,-5)and(-5,\infty)\]

OpenStudy (anonymous):

Are they above or below the x-axis?

OpenStudy (anonymous):

Above the x axis.

OpenStudy (anonymous):

Both?

OpenStudy (anonymous):

yeah. Graph it and see.

OpenStudy (anonymous):

The vertex touches the x axis at the point (-5,0) but never goes below it.

OpenStudy (anonymous):

Yes thus its an even multiplicity

OpenStudy (anonymous):

Thanks